The largest swimming pool in the world is in Casablanca. With a perimeter of 1110 meters this rectangular shaped pool is 30 meters longer than 6 times its width. How many meters wide is it?

Answer by Alan3354(69443)

The largest swimming pool in the world is in Casablanca. With a perimeter of 1110 meters this rectangular shaped pool is 30 meters longer than 6 times its width. How many meters wide is it?
---------------
2W + 2L = 1110
W + L = 550
---------
L = 6W + 30
--
W + 6W + 30 = 550
7W = 520
W = 520/7 meters
==================
http://www.victory-cruises.com/pool.html
The pool in Chile is a LOT bigger.